OpenStudy (anonymous):

The function f(x) = 4(3)^x represents the growth of a dragonfly population every year in a remote swamp. Erin wants to manipulate the formula to an equivalent form that calculates four times a year, not just once a year. Which function is correct for Erin's purpose, and what is the new growth rate? f(x) = 4(3)x; growth rate 300% f(x) = 4(3)x, growth rate 4% f(x) = 4(1.32)x; growth rate 4% f(x) = 4(1.32)4x; growth rate 32%

OpenStudy (anonymous):

f(x) = 4(3)^x ?

OpenStudy (anonymous):

actually there is a right answer here.. do you know the compound interest formula?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

yeah, you need that

OpenStudy (irishboy123):

to solve this mechanically, just re-write \(f(x) = 4 (3^x)\) as \(f(x) = 4 (3^{4x})^{1/4} = 4 (3^{1/4})^{4x}\) then calculate \(3^{1/4}\) however, i am totally unsure that any of this makes any practical sense